OpenStudy (anonymous):

Simplify into a + bi. -9+8i/ 1 +2i

mathslover (mathslover):

\[\large{\frac{-9+8i}{1+2i}}\]

Parth (parthkohli):

Oh, don't you love making new accounts and posting the same question again and again?

Parth (parthkohli):

Well, do you know about conjugates?

OpenStudy (anonymous):

Well everyone was just bring rude, so I figured I would try again and see if someone serious would really help me.

mathslover (mathslover):

No problem, fortunately , I am not going to be rude, unless you start making no effort, answer the question of parth.

OpenStudy (anonymous):

Thank you mathslover for helping. I didn't mean you. :)

mathslover (mathslover):

:) , though, I didn't help you yet, Do you know about conjugates?

OpenStudy (anonymous):

I am re-learning them. It has been a while. I have forgotten.

mathslover (mathslover):

ok , so an introduction to conjugates (you need) : \(\large{\color{blue}{\textbf{conjugate of a+bi will be :}} = \color{red}{a-bi}}\) that is the + sin became - . Getting it?

OpenStudy (anonymous):

Yes. So -9-8i/1-2i?

mathslover (mathslover):

No, see: \(\large{\frac{a}{b} = \frac{ab}{b\times b} = \frac{ab}{b^2}}\) right?

OpenStudy (anonymous):

I suppose.

mathslover (mathslover):

Right! so , \(\large{\color{blue}{\frac{-9+8i}{1+2i}}=\color{blue}{\frac{-9+8i}{1+2i}}\times \color{green}{\frac{1-2i}{1-2i}}}\) What I did is : I just multiplied numerator and denominator with the conjugate of denominator i.e conjugate of 1+2i .

mathslover (mathslover):

Can you tell me what is the conjugate of 1+2i ?

OpenStudy (anonymous):

1-2i

mathslover (mathslover):

good so that is what I multiplied. Getting it?

OpenStudy (anonymous):

Yes.

mathslover (mathslover):

So , can you tell me what will be : \[\huge{\color{blue}{\frac{(-9-8i)(1-2i)}{(1+2i)(1-2i)}}}\]

mathslover (mathslover):

oops there is -9+8i in the numerator.

OpenStudy (anonymous):

Is that correct? Can I proceed?

OpenStudy (anonymous):

-9-16i/1-2i.....Would that be correct?

mathslover (mathslover):

\[\LARGE{\frac{(-9+8i)(1-2i)}{(1+2i)(1-2i)}}\] First solve for this!

OpenStudy (anonymous):

I did. :) Well I tried...

mathslover (mathslover):

What you got in numerator?

OpenStudy (anonymous):

I got -9-16i/-1-2i

mathslover (mathslover):

I think you did mistake somewhere, well let us solve it!

mathslover (mathslover):

(-9+8i)(1-2i) ^Let us first solve for numerator, OK? Ready?

OpenStudy (anonymous):

Sorry. It is in there somewhere, I am just refreshing my memory. Haven't done these since high school.

OpenStudy (anonymous):

Yes.

mathslover (mathslover):

No problem, I am here to refresh dude, cool down. OK, so : can you expand this : (-9+8i)(1-2i) ?

mathslover (mathslover):

just like you do : (a+b)(c+d) = ac + ad + bc + bd Expand (-9+8i)(1-2i) .

OpenStudy (anonymous):

I don't get that.

mathslover (mathslover):

ok , expand (2 + a)(3+b)

OpenStudy (anonymous):

2 + a * 3 + b

mathslover (mathslover):

See: use distributive property to expand : (2+a)(3+b) 2(3) + 2(b) +a(3) +a(b) = 6+2b+3a+ab Getting it?

OpenStudy (anonymous):

Slightly. It is kind of confusing.

mathslover (mathslover):

hmn, do you know about distributive property?